### v5

@@ 4.2. Benchmarks: Comparative Assessment
This section formalizes our governance-aligned rubric (C1–C10) and applies it to widely used benchmarks; we then compare coverage and link results to audit evidence in the Tables 6 and 8.
@@ 4.2.1. Comparative Assessment Protocol
We assess widely used AI safety benchmarks against a ten–item rubric aligned to risk families and governance needs: C1 Bias/Discrimination, C2 Toxicity, C3 Security/Robustness (including adversarial), C4 Misinformation/Disinformation, C5 Deepfakes/Media Integrity, C6 Privacy (membership/extraction), C7 System–level failure, C8 Malicious–actor realism (red–team/adaptive), C9 Metric validity/documentation, and C10 Governance alignment (e.g., NIST AI RMF, EU AI Act). The ten criteria (also given in Appendix B) reflect risk families that recur across governance instruments and risk repositories, including the NIST AI Risk Management Framework, the EU AI Act, and ISO/IEC 42001, as well as consolidated mappings such as the MIT AI Risk Repository (nis, 2024; Outeda, 2024; Systems, 2021; Slattery et al., 2024). We scored each benchmark after reviewing the official paper, public documentation, and widely cited usage reports; disagreements were resolved by discussion and, when needed, a third reviewer.
Each benchmark is scored $s_{i}(B)\in\{0,1,2\}$ per criterion (0 = no coverage, 1 = partial, 2 = strong). Equal weights are used by default. The normalized score is
$S(B)=\frac{\sum_{i=1}^{10}w_{i}\,s_{i}(B)}{2\sum_{i=1}^{10}w_{i}},\quad w_{i}>0.$
This yields values in $[0,1]$, where 0 indicates no coverage and 1 indicates full coverage across all criteria. We also provide a governance crosswalk indicating which requirements each benchmark can plausibly evidence (Tables 6 and 8).
4.2.2. Comparative Assessment Findings
Building on the challenges in Section 3.2, we apply the rubric in Section 4.2.1 to evaluate benchmarks against governance needs. Table 5 presents the risk surface matrix for criteria C1–C8. Criteria C9 (metric validity and documentation) and C10 (governance alignment) are summarized in Table 8.
@@ 4.2.2. Comparative Assessment Findings
Table 5. Mapping benchmarks to risk surface areas. ✓ = supported; $\triangle$ = partial; – = no explicit coverage (from public papers and documentation).
@@
Benchmark Bias/Discr. Toxicity Security Misinfo/Disinfo Deepfakes Privacy System failure Malicious actors
HarmBench (Mazeika et al., 2024) $\triangle$
SALAD Bench (Li et al., 2024a) $\triangle$ $\triangle$ $\triangle$
ALERT (Tedeschi et al., 2024) $\triangle$ $\triangle$
Risk Taxonomy (Cui et al., 2024) $\triangle$ $\triangle$ $\triangle$
TrustGPT (Huang et al., 2023b)
RealToxicityPrompts (Gehman et al., 2020)
DecodingTrust (Wang et al., 2023a) $\triangle$ $\triangle$ $\triangle$
SafetyBench (Zhang et al., 2024) $\triangle$ $\triangle$ $\triangle$
MM–SafetyBench (Liu et al., 2024) $\triangle$ $\triangle$ $\triangle$ $\triangle$
XSTest (Röttger et al., 2024) $\triangle$
Rainbow Teaming (Samvelyan et al., 2024) $\triangle$ $\triangle$
MLCommons AI Safety v0.5 (Vidgen and et al., 2024) $\triangle$ $\triangle$ $\triangle$
HELM (Liang et al., 2023) $\triangle$ $\triangle$
SHIELD (Shi et al., 2024)
BIG-bench (bench authors, 2023) $\triangle$ $\triangle$ $\triangle$ $\triangle$
BEADs (Raza et al., 2024b) $\triangle$ $\triangle$
PrivLM-Bench (Li et al., 2024c) $\triangle$ $\triangle$
PrivacyLens (Shao and et al., 2024)
HumaniBench (Raza et al., 2025a) $\triangle$ $\triangle$ $\triangle$
@@
Table 6. Comparative assessment rubric for AI safety benchmarks (0 = no coverage, 1 = partial, 2 = strong). Scores use equal weights and are normalized to $[0,1]$.
@@
HumaniBench (Raza et al., 2025a) 2 1 0 1 0 1 1 0 1 1 0.40
Criteria (C1 to C10). C1 Bias/Discrimination; C2 Toxicity; C3 Security/Robustness; C4 Misinformation/Disinformation; C5 Deepfakes/Media Integrity; C6 Privacy (membership/extraction); C7 System level failure (tool/agentic); C8 Malicious actor realism (red team/adaptive); C9 Metric validity/documentation; C10 Governance alignment. Normalization (equal weights): $S(B)=\frac{\sum_{i=1}^{10}s_{i}(B)}{20}$ with $s_{i}\in\{0,1,2\}$.
Table 7. Governance-based Frameworks and Technical AI Tools used in RAI Practices.
@@
Fairness & Bias Mitigation ISO/IEC TR 24027:2021 (3, 2021) : technical report on AI and data quality. NIST SP-1270 (Schwartz et al., 2022) : framework for identifying and managing bias in AI. Framework for increasing diversity & inclusion in AI education, research, & policy: AI4ALL (AI4ALL, 2024) Evaluate for model fairness via visualizations: Google What-if Tool (Wexler et al., 2020) AI Fairness 360 Toolkit by IBM provides tools for evaluation and mitigation of bias: AIF360 (Bellamy et al., 2018b) Assess and enhance fairness of ML models: Microsoft Fairlearn (Weerts et al., 2023)
@@
Interpretability & Explainability IEEE P2976 (Polemi et al., 2024) : Standards for Explainable Artificial Intelligence. IEEE P2894 (Committee, 2024) : guidelines for implementing explainable AI technologies. Explaining Decisions Made with AI: UK Info. Commissioner’s Office & The Alan Turning Institute (Leslie, 2022) Model interpretability and explainability tool: LIT (Tenney et al., 2020) Toolkit for measuring explainability using various models like LIME and SHAP: InterpretML (Sameki et al., 2020) Provide simple AI generated explanations: ELI5 (Deepgram, 2024) AI Explainability 360 toolkit by IBM provides tools for explainability of ML models: AIX360 (Arya et al., 2019)
Transparency & Accountability U.S Government Accountability Office: AI Accountability Framework (U.S. Government Accountability Office, 2021) Organisation for Economic Co-operation and Development (OECD): AI Principle of Accountability ((, OECD) Evaluate, test and monitor ML models: Evidently (EvidentlyAI, 2024) Improve model transparency, reproducibility, and robustness: MLBench (MLBench, 2020)
@@
Privacy, Safety & Security NASA (Smith et al., 2020): hazard contribution modes of ML in space applications. IEEE P7009 (Farrell et al., 2021): standards for the fail-safe design of autonomous systems. ISO/IEC TS 27022:2021 (Systems, 2021): information security management systems guidelines. NIST AI 100-2 E2023 (Oprea and Vassilev, 2023) : specific cybersecurity practices for AI. UK Government AI Cybersecurity Code of Practice (Department of Science, Innovation & Technology, 2024). MIT Risk repo (Slattery et al., 2024): Database documenting AI risks and safety gaps. Google’s Secure AI Framework SAIF (Inc., 2024) Context-aware toxicity tracker: Unitary (Unitary, 2023) Adversarial Robustness Toolbox: ART Audit privacy of ML models: Privacy Meter Google Differential Privacy libraries: DP Privacy-preserving data analysis and ML toolkit Secret Flow Tensorflow privacy (TensorFlow, 2023) Privacy Enhancing Technology (PET) on Synthetic Data Generation: Betterdata.AI (inc., 2024)
Ethical Guidelines & Compliance European General Data Protection Regulation: GDPR (Union, 2018) Canadian AI and Data Act: ISED (of Canada, 2023a) Canadian Voluntary Code of Conduct on the Responsible Development and Management of Advanced GenAI Systems: ISED (of Canada, 2023b) American Executive Order on Safe, Secure, and Trustworthy Artificial Intelligence: AEO (House, 2023) Google responsible AI practices (AI, 2023) Australia’s AI Ethics Principles (Australian Government Department of Industry, Science and Resources, 2019) Microsoft’s Responsible AI Toolbox (Microsoft, 2024) Demonstrate trust, measure and manage risk, and compliance: OneTrust (OneTrust, 2023) World Health Organization (WHO): Ethics and governance of AI for health Guidance on large multi-modal models (Organization et al., 2024)
@@
Monitoring & Auditing Index of harms realized in the real world by AI systems: Incident Database (Database, 2023) Guidance provided by UK Information Commissioner’s Office AI Auditing Framework (UK Government Information Commissioner’s Office, 2020) Audit classification models for intersectional bias: FairVis (Cabrera et al., 2019) Bias and fairness audit toolkit: Aequitas (for Data Science and of Chicago, 2018) Auditing of bias in generalized ML applications: Audit AI (Inc., 2020)
Table 8. Governance alignment crosswalk (illustrative). ✓= supported evidence; $\triangle$= partial; – = not covered.
@@
NIST: Privacy & Security (nis, 2020a, b) $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$
@@
Multimodal Deepfake Risk (c2p, 2024) $\triangle$
Adaptive/Red-team Evidence (nis, 2023) $\triangle$ $\triangle$ $\triangle$ $\triangle$
@@
Abbreviations. RTP = RealToxicityPrompts; MM–Saf. = MM–SafetyBench; HarmB. = HarmBench; MLC v0.5 = MLCommons AI Safety v0.5; Humani = HumaniBench.
@@
Risk-surface coverage. Table 5 presents the representative benchmarks cover eight salient risk families (bias/discrimination, toxicity, security/robustness, mis/disinformation, deepfakes/media integrity, privacy, system-level failure, and malicious-actor realism). Entries are based on public papers and documentation and reflect explicit task/metric coverage rather than incidental behavior. The results in Table 5 shows that coverage is densest for toxicity and bias (e.g., RealToxicityPrompts, SafetyBench, HarmBench), while privacy is siloed to specialized suites (PrivLM-Bench, PrivacyLens). Deepfake/media integrity is sparsely addressed (notably SHIELD), and only a few efforts incorporate adaptive adversaries or richer system context (Rainbow Teaming, ALERT). System-failure probing is rare, appearing primarily in Rainbow Teaming.
@@
Rubric-scored comparison. We operationalize the ten-criterion rubric in Section 4.2.1 to produce a normalized [0,1] score per benchmark. The aim is not to rank “best overall”, but to indicate breadth and governance-ready evidence across risk families and evaluation qualities (metric validity, documentation, policy alignment). The results in Table 6 shows that Rainbow Teaming attains the highest breadth score (0.50), reflecting strong coverage of security under adaptive attack and red-team realism. Risk Taxonomy and HumaniBench (0.40 each) combine multi-risk awareness with governance-facing structure. MLCommons AI Safety v0.5 and DecodingTrust (0.35) provide broad but mostly static probes. Single-issue suites score lower by design,e.g., RealToxicityPrompts and PrivacyLens (0.15), useful for their focus but limited for governance evidence.
@@
Governance context and crosswalk. Table 7 situates the analysis within widely used frameworks and tools. Table 8 then maps each benchmark to selected governance requirements spanning NIST AI RMF privacy/security controls, EU AI Act robustness and transparency obligations, multimodal deepfake risk, and evidence for adaptive/red-team evaluation. A ✓indicates plausible evidentiary use; $\triangle$ indicates partial or indirect support; indicates no explicit coverage in public artifacts. The results show that only a subset of suites provide tangible evidence for privacy/security controls (e.g., PrivLM-Bench). Robustness/resilience support is strongest where adaptive red teaming is integral (Rainbow Teaming, ALERT). Deepfake risk is narrowly covered (SHIELD), suggesting a gap for multimodal governance. Transparency/record-keeping signals are uneven outside of meta-benchmarks (HELM, MLCommons), which can hinder audit trails under the EU AI Act.

### v6

@@ 4.2. Benchmarks: Comparative Assessment
This section formalizes our governance-aligned rubric (C1–C10) and applies it to widely used benchmarks; we then compare coverage and link results to audit evidence in Tables 4 and 6.
@@ 4.2.1. Comparative Assessment Protocol
Our benchmark assessment follows a structured qualitative expert appraisal with ordinal scoring, a well-established approach for rubric-based evaluation in computer science survey research (Harzing and Alakangas, 2016). It is neither a quantitative meta-analysis (no effect sizes or pooled statistics are computed) nor an unstructured qualitative review. Instead, we employ codebook-guided independent scoring that yields ordinal coverage indicators (0/1/2 per criterion), which are then normalized into a composite index. The resulting scores characterize the breadth of risk-surface coverage and documentation quality across benchmarks (they are not cardinal performance measures and should not be interpreted as rankings of overall safety). This approach is appropriate because the underlying evidence, e.g., benchmark papers, dataset cards, public documentation, and is heterogeneous in format and scope, precluding formal quantitative synthesis.
Rubric design.
@@
We assess widely used AI safety benchmarks against a ten-item rubric aligned to risk families and governance needs: C1: Bias/Discrimination, C2 :Toxicity, C3 : Security/Robustness (including adversarial), C4: Misinformation/Disinformation, C5 : Deepfakes/Media Integrity, C6 : Privacy (membership/extraction), C7 : System-level failure, C8 : Malicious-actor realism (red-team/adaptive), C9 : Metric validity/documentation, and C10: Governance alignment (e.g., NIST AI RMF, EU AI Act). The ten criteria (also given in Appendix A.2) reflect risk families that recur across governance instruments and AI risk repositories.
Scoring procedure and reliability.
Each benchmark was scored $s_{i}(B)\in\{0,1,2\}$ per criterion (0 = no coverage, $1$ = partial, $2$ = strong) by two independent reviewers, following the codebook in Appendix A.2. The codebook specifies per-criterion construct definitions, qualifying evidence types, scoring anchors, and a disambiguation rule that assigns tasks spanning multiple risk families to the primary stated objective to prevent double-counting. Evidence was drawn according to a priority ordering: (i) official paper and documentation; (ii) dataset or model cards and maintained leaderboards; (iii) widely cited third-party reports. Disagreements were resolved through structured discussion and, where necessary, adjudication by a third reviewer. Equal weights are used by default. The normalised score is
(1) $S(B)=\frac{\sum_{i=1}^{10}w_{i}\,s_{i}(B)}{2\,\sum_{i=1}^{10}w_{i}},\quad w_{i}>0.$
This yields values in $[0,1]$, where 0 indicates no coverage and $1$ indicates full coverage across all criteria. To verify that our conclusions are not an artefact of the equal-weight assumption, we conducted a sensitivity analysis under alternate weight profiles that prioritise high-stakes criteria (e.g., C3 Security, C6 Privacy); the resulting Spearman rank correlation exceeded $\rho>0.85$, indicating that comparative breadth is a stable indicator of evaluative maturity.
@@
Table 3. Mapping benchmarks to risk surface areas. $\checkmark$ = supported; $\triangle$ = partial; – = no explicit coverage (from public papers and documentation). This crosswalk expands on rubric criteria C9 (metric validity/documentation) and C10 (governance alignment) from Table 4.
@@
Benchmark Bias/Discrimination Toxicity Security Misinfo Deepfakes Privacy System fail. Malicious
HarmBench (Mazeika et al., 2024) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
SALAD-Bench (Li et al., 2024a) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\triangle$
ALERT (Tedeschi et al., 2024) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Risk Taxonomy (Cui et al., 2024) $\checkmark$ $\checkmark$ $\triangle$ $\triangle$ $\triangle$ $\triangle$
TrustGPT (Huang et al., 2023b) $\checkmark$ $\checkmark$
RealToxicityPrompts (Gehman et al., 2020) $\checkmark$
DecodingTrust (Wang et al., 2023a) $\checkmark$ $\checkmark$ $\checkmark$ $\triangle$ $\triangle$ $\triangle$
SafetyBench (Zhang et al., 2024) $\checkmark$ $\checkmark$ $\triangle$
MM-SafetyBench (Liu et al., 2024) $\checkmark$ $\triangle$ $\triangle$ $\triangle$ $\checkmark$
XSTest (Röttger et al., 2024) $\triangle$ $\checkmark$ $\triangle$
Rainbow Teaming (Samvelyan et al., 2024) $\triangle$ $\checkmark$ $\triangle$ $\checkmark$ $\checkmark$
MLCommons AI Safety v0.5 (Vidgen and et al., 2024) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
HELM (Liang et al., 2023) $\checkmark$ $\checkmark$ $\triangle$ $\checkmark$
SHIELD(Shi et al., 2024) $\triangle$ $\triangle$ $\checkmark$
BIG-bench (bench authors, 2023) $\checkmark$ $\checkmark$ $\triangle$ $\triangle$ $\triangle$
PrivLM-Bench (Li et al., 2024b) $\checkmark$
PrivacyLens (Shao and et al., 2024) $\triangle$ $\checkmark$
HumaniBench (Raza et al., 2025) $\checkmark$ $\triangle$ $\triangle$ $\triangle$ $\triangle$
@@
Table 4. Comparative assessment rubric for AI safety benchmarks (0 = no coverage, 1 = partial, 2 = strong). Scores use equal weights and are normalized to $[0,1]$. Per-criterion scoring anchors and evidence rules are provided in Appendix A.2. This table reports scores across all ten criteria (C1–C10), including C9 (metric validity/documentation) and C10 (governance alignment).
@@
HumaniBench (Raza et al., 2025) 2 1 0 1 0 1 1 0 1 1 0.40
Criteria (C1 to C10) C1 Bias/Discrimination; C2 Toxicity; C3 Security/Robustness; C4 Misinformation/Disinformation; C5 Deepfakes/Media Integrity; C6 Privacy (membership/extraction); C7 System level failure (tool/agentic); C8 Malicious actor realism (red team/adaptive); C9 Metric validity/documentation; C10 Governance alignment. Normalization (equal weights): $S(B)=\frac{\sum_{i=1}^{10}s_{i}(B)}{20}$ with $s_{i}\in\{0,1,2\}$.
Table 5. Governance-based Frameworks and Technical AI Tools used in RAI Practices
@@
Fairness & Bias Mitigation ISO/IEC TR 24027:2021 (ISO/IEC, 2021a): technical report on AI and data quality. NIST SP-1270 (Schwartz et al., 2022): framework for identifying and managing bias in AI. AI4ALL (AI4ALL, [n. d.]): framework for increasing diversity & inclusion in AI education, research, & policy. African Union (AU) Data Policy Framework (African Union, 2022): promotes equitable, inclusive AI and data governance across member states. Singapore IMDA & PDPC’s AI Governance Framework (Infocomm Media Development Authority and Personal Data Protection Commission (PDPC)(2020), IMDA): addresses fairness, human-centricity, and non-discrimination in AI decision-making. Google What-If Tool (Wexler et al., 2020): evaluate for model fairness via visualizations. IBM AI Fairness 360 (AIF360) (Bellamy et al., 2019): tools for evaluation and mitigation of bias. Microsoft Fairlearn (Weerts et al., 2023): assess and enhance fairness of ML models.
@@
Interpretability & Explainability IEEE P2976 (Polemi et al., 2024): standards for Explainable Artificial Intelligence. IEEE P2894 (Committee, 2024): guidelines for implementing explainable AI technologies. UK ICO & The Alan Turing Institute: Explaining Decisions Made with AI (Information Commissioner’s Office and The Alan Turing Institute, 2022). China CAC Algorithm Recommendation Provisions (Cyberspace Administration of China, 2022b): mandates transparency and explainability requirements for algorithmic recommendation systems. Learning Interpretability Tool (LIT) (Google PAIR, [n. d.]): model interpretability and explainability. InterpretML (Microsoft Research): toolkit for measuring explainability using LIME and SHAP. IBM AI Explainability 360 (AIX360) (Arya et al., 2019): tools for explainability of ML models.
Transparency & Accountability U.S. Government Accountability Office: AI Accountability Framework (U.S. Government Accountability Office, 2021). OECD AI Principle of Accountability (Organisation for Economic Co-operation and Development, 2019). China Measures for the Management of Generative AI Services (CAC, 2023) (Cyberspace Administration of China, 2023): requires service providers to disclose AI-generated content, maintain logs, and be accountable for model outputs. Brazil’s Proposed AI Act (PL 2338/2023) (Senado Federal do Brasil, 2023): mandates impact assessments, transparency obligations, and accountability measures for high-risk AI systems. SDAIA National AI Ethics Principles (Saudi Data and Artificial Intelligence Authority, 2023a): establishes accountability, auditability, and transparency requirements for AI deployments in Saudi Arabia. Evidently (Evidently AI, [n. d.]): evaluate, test and monitor ML models. MLBench (MLBench Team, [n. d.]): improve model transparency, reproducibility, and robustness. AI Verify (Infocomm Media Development Authority ([n. d.]), IMDA): Singapore’s IMDA-developed AI governance testing framework.
@@
Privacy, Safety & Security NASA (Smith and Kearns, 2020): hazard contribution modes of ML in space applications. IEEE P7009 (Farrell et al., 2021): standards for the fail-safe design of autonomous systems. ISO/IEC TS 27022:2021 (ISO/IEC, 2021b): information security management systems guidelines. NIST AI 100-2 E2023 (Oprea and Vassilev, 2023b): specific cybersecurity practices for AI. UK Government AI Cybersecurity Code of Practice (Department of Science, 2024). MIT AI Risk Repository (Slattery et al., 2024): database documenting AI risks and safety gaps. Google’s Secure AI Framework (SAIF) (Google, 2023). Brazil LGPD (da Republica, Brasil): data protection law governing privacy, data processing, and individual rights in AI-driven systems. Nigeria Data Protection Act (NDPA, 2023) (Agency, 2023): establishes data privacy safeguards for AI systems processing personal data of Nigerian citizens. China Personal Information Protection Law (of China, 2021): comprehensive data privacy regulation for AI systems. UAE Federal Decree-Law No. 45 on Personal Data Protection (Government, 2021): governs data privacy and individual rights in AI-driven data processing within the UAE. Unitary (Ltd., [n. d.]): context-aware toxicity tracker. Adversarial Robustness Toolbox (ART) (Research, [n. d.]). Privacy Meter (Privacy and Lab, [n. d.]): audit privacy of ML models. Google Differential Privacy Libraries (Google, [n. d.]). SecretFlow (Group, [n. d.]): privacy-preserving data analysis and ML toolkit. TensorFlow Privacy (Google, [n. d.]). Betterdata.AI (inc., 2024): PET on synthetic data generation.
Ethical Guidelines & Compliance European General Data Protection Regulation (GDPR) (European Parliament and Council of the European Union, 2016). Canadian AI and Data Act (Innovation, Science and Economic Development Canada, 2023a). Canadian Voluntary Code of Conduct on the Responsible Development and Management of Advanced GenAI Systems (Innovation, Science and Economic Development Canada, 2023b). American Executive Order on Safe, Secure, and Trustworthy AI (House, 2023). China’s New Generation AI Development Plan (AIDP, 2017) & AI Governance Principles (2019) (State Council of the People’s Republic of China, 2017; National New Generation AI Governance Expert Committee (2019), China): outline China’s national strategy for AI development with ethical principles for security, fairness, and human-centric AI. UAE National AI Strategy 2031 (United Arab Emirates Government, [n. d.]): comprehensive national strategy establishing ethical AI goals and responsible AI deployment. SDAIA Responsible AI Guidelines (Saudi Data and Artificial Intelligence Authority, 2023b): ethical guidelines for responsible AI use and regulatory compliance in Saudi Arabia. Singapore Model AI Governance Framework (2nd Ed., IMDA) (Infocomm Media Development Authority and Personal Data Protection Commission (PDPC)(2020), IMDA): practical guidance on responsible AI deployment covering internal governance and human oversight. AU Continental AI Policy Framework (2024) (African Union, 2024): continent-wide framework for ethical, inclusive, and human rights-respecting AI governance. Colombia’s AI Ethics Framework (Ministerio de Tecnologías de la Información y las Comunicaciones (2019), MinTIC) & Chile National AI Policy (Ministerio de Ciencia, Tecnología, Conocimiento e Innovación, 2021): regional strategies for responsible AI adoption. Google Responsible AI Practices (Google, [n. d.]). Australia’s AI Ethics Principles (Australian Government Department of Industry, Science and Resources, [n. d.]). Microsoft’s Responsible AI Toolbox (principles and approach, 2024). OneTrust (OneTrust LLC, [n. d.]): demonstrate trust, measure and manage risk, and compliance. WHO: Ethics and Governance of AI for Health (World Health Organization, 2024). Project Moonshot (AI Verify Foundation, 2024): open-source LLM evaluation and red-teaming toolkit to benchmark GenAI systems against safety and ethical compliance standards.
@@
Monitoring & Auditing AI Incident Database (Responsible AI Collaborative, [n. d.]): index of harms realized in the real world by AI systems. UK ICO AI Auditing Framework (UK Information Commissioner’s Office (2020), ICO). China CAC Measures for the Management of Deep Synthesis Internet Information Services (2022) (Cyberspace Administration of China, 2022a): requires labeling, monitoring, and auditing of AI-generated synthetic media. Smart Africa Alliance AI Audit Guidelines (Smart Africa, 2023): monitoring and audit guidance for AI systems in African public services. IDB AI Governance Toolkit (Inter-American Development Bank (2023), IDB): monitoring and assessment tools for responsible AI deployment across Latin American and Caribbean nations. FairVis (Poloclub, [n. d.]): audit classification models for intersectional bias. Aequitas (Center for Data Science and Public Policy, [n. d.]): bias and fairness audit toolkit. Audit-AI : auditing of bias in generalized ML applications.
Table 6. Governance alignment crosswalk (illustrative). ✓= supported evidence; $\triangle$= partial; –= not covered. A ✓ indicates plausible evidentiary support rather than a statement of legal compliance; see Appendix B for the non-compliance note. This crosswalk expands on rubric criteria C9 (metric validity/documentation) and C10 (governance alignment) from Table 4.
@@
NIST: Privacy & Security (of Standards and, NIST, NIST) $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$
@@
Multimodal Deepfake Risk (C2PA, 2024) $\triangle$
Adaptive/Red-team Evidence (of Standards and, NIST) $\triangle$ $\triangle$ $\triangle$ $\triangle$
@@
Abbreviations RTP = RealToxicityPrompts; MM–Saf. = MM–SafetyBench; HarmB. = HarmBench; MLC v0.5 = MLCommons AI Safety v0.5; Humani = HumaniBench.
4.2.2. Comparative Assessment Findings.
Building on the challenges in Section 3.2, we apply the rubric in Section 4.2.1 to evaluate benchmarks against governance needs. Table 3 presents the risk-surface matrix for criteria C1–C8. The full rubric scores, including C9 (metric validity and documentation) and C10 (governance alignment), are given in Table 4. Table 6 then provides a finer-grained governance crosswalk mapping each benchmark to specific regulatory requirements.
Risk-surface coverage.
@@
Table 3 presents the representative benchmarks across eight salient risk families (C1–C8). Entries are based on public papers and documentation and reflect explicit task/metric coverage rather than incidental behaviour. The results in Table 3 show that coverage is densest for toxicity and bias (e.g., RealToxicityPrompts, SafetyBench, HarmBench), while privacy is siloed to specialized suites (PrivLM-Bench, PrivacyLens). Deepfake/media integrity is sparsely addressed (notably SHIELD), and only a few efforts incorporate adaptive adversaries or richer system context (Rainbow Teaming, ALERT). System-failure probing is rare, appearing primarily in Rainbow Teaming.
Rubric-scored comparison.
@@
We operationalize the ten-criterion rubric in Section 4.2.1 to produce a normalized $[0,1]$ score per benchmark. The aim is not to rank “best overall,” but to indicate breadth and governance-ready evidence across risk families and evaluation qualities (metric validity, documentation, policy alignment). The results in Table 4 show that Rainbow Teaming attains the highest breadth score (0.50), reflecting strong coverage of security under adaptive attack and red-team realism. Risk Taxonomy and HumaniBench (0.40 each) combine multi-risk awareness with governance-facing structure. MLCommons AI Safety v0.5 and DecodingTrust (0.35) provide broad but mostly static probes. Single-issue suites score lower by design, e.g., RealToxicityPrompts and PrivacyLens (0.15), useful for their focus but limited for governance evidence.
Governance context and crosswalk.
@@
Table 5 situates the analysis within widely used frameworks and tools. Table 6 then maps each benchmark to selected governance requirements and risks such as multimodal deepfake risk, and evidence for adaptive/red-team evaluation. The results show that only a subset of suites provide tangible evidence for privacy/security controls (e.g., PrivLM-Bench). Robustness/resilience support is strongest where adaptive red teaming is integral (Rainbow Teaming, ALERT). Deepfake risk is narrowly covered (SHIELD), suggesting a gap for multimodal governance. Transparency/record-keeping signals are uneven outside of meta-benchmarks (HELM, MLCommons), which can hinder audit trails under the EU AI Act.
Figure 8. Responsible GenAI governance landscape and linkage to our evaluation criteria. Technical factors correspond primarily to rubric items C1–C7 (fairness/bias, toxicity, security/robustness, mis/disinformation, deepfakes/media integrity, privacy, system-level failure). Risk & assurance processes (anticipation, responsiveness, reflexivity, inclusion, knowledge management) support adaptive testing (C8), metric validity/record-keeping (C9) and Legal instruments (e.g., EU AI Act, NIST AI RMF) ground C10 (governance alignment). Sustainable AI highlights environmental/societal/economic impacts that are only weakly covered by current benchmarks, motivating the gaps observed in Tables 46